# RLHFuse: Efficient RLHF Training for Large Language Models with Inter- and Intra-Stage Fusion

### v2

@@ 7.3 Effectiveness of Stage Fusion
Inter-stage fusion. One key parameter in inter-stage fusion is the migration threshold $R_{t}$. As $R_{t}$ increases from zero, it shifts from fully serial execution to more aggressive fused execution. However, if $R_{t}$ is set too large, the long-tail generation instances may be overloaded, extending the original generation time. We measure the execution time of fused generation and inference stage under different migration ratios ($R_{t}/BS$) as shown in Figure 9. We can see the optimal latency is achieved when the remaining samples account for about 20% of the batch size. Note that $R_{t}$ is related to the output length distribution of the model and this experiment is conduct during the early stage of training. As training progresses, $R_{t}$ requires periodic adjustments to adapt the distribution change. In practice, such changes are usually not too drastic, so the adjustment frequency remains low.
Settings Latency Speedup Memory Usage
@@
Models PP0 PP1 #MB greedy ours optimal greedy ours
33B/13B 8 4 8 1.29 1.38 1.38 1.51 1.0
16 1.12 1.30 1.30 2.14 1.0
32 1.10 1.15 1.15 2.75 1.26
8 8 1.4 1.4 1.4 1.51 1.0
16 1.15 1.32 1.32 2.00 1.19
32 1.08 1.17 1.17 2.75 1.31
65B/33B 16 8 16 1.28 1.48 1.48 1.61 1.0
32 1.14 1.27 1.27 2.23 1.0
64 1.12 1.15 1.15 2.88 1.26
16 16 1.27 1.5 1.5 1.57 1.0
32 1.16 1.33 1.33 2.00 1.22
64 1.09 1.16 1.17 2.88 1.47
Table 3: The latency speedup and peak activation memory cost of pipeline schedules found by different algorithms under different model, pipeline stage, and micro-batch settings.
Intra-stage fusion. We compare RLHFuse’s simulated annealing algorithm against the greedy approach mentioned in §5.2 and show the latency speedup and peak activation memory usage relative to the serial execution of the two models with 1F1B schedule in Table 3. For the latency, we also include a lower bound calculated as follows: we first compute the earliest possible completion time on each stage, which consists of three parts: the earliest possible arrival time of the first task, the total time required to process all the tasks, and the remaining time needed for the subsequent pipeline stages of the final task; we then take the maximum of these times across all stages as the lower bound of the fused pipeline schedule. Note that there may not necessarily exist a schedule that reaches this lower bound, but we can use it as a metric to assess how close our algorithm is to the optimal solution. Under all the settings, our approach outperforms the greedy method and achieves the theoretical lower bound except for the last case. As for the memory cost, the serial 1F1B execution serves as the lower bound for the fused schedule. Our method also greatly outperforms the greedy approach and often reaches the lower bound. Even when it does not achieve the lower bound, the extra overhead remains within an acceptable range.
@@ 7.4 Deep Dive
In this section, we deep dive into RLHFuse by showing a fused pipeline schedule produced by RLHFuse for the 65B/33B model setting. The upper of Figure 10 shows the GPU execution timeline, with forward and backward tasks in light and dark colors, respectively. With intra-stage fusion, RLHFuse strategically fuses one 65B model of 16 PP stages (blue grids) with two 33B models that each has 8 PP stages (pink grids) to fill the pipeline bubbles of each other in reversed pipeline directions. Ultimately, the execution time of this fused pipeline schedule is the same as individually executing the 65B model using 1F1B schedule, which means we completely overlap the training of the 33B model, achieving the theoretical lower bound. Meanwhile, as shown in the bottom of Figure 10, the peak activation memory usage also achieves the lower bound of executing the two models in serial with 1F1B schedule. This complex schedule, which achieves optimal performance in both time and memory usage, is difficult to design manually, further demonstrating the effectiveness of our fused pipeline schedule generation algorithm proposed in §5.2.
@@ 8 Related Work
RLHF training systems. As RLHF gradually becomes the primary approach for LLM alignment, many distributed frameworks are specifically designed for RLHF training. trlX [46] proposes algorithmic system enhancements to reduce compute and memory requirements. OpenRLHF [47] notices that different RLHF training tasks have different resource requirements and proposes to designate customized parallel strategies for different tasks. ReaLHF [13] further proposes parameter reallocation to dynamically redistribute LLM parameters on shared GPU devices to avoid under-utilization. PUZZLE [14] proposes lightweight context switching to reduce the task switching overhead, which is orthogonal and partly integrated into RLHFuse. However, all the above systems overlook the subtask-level optimization opportunities and suffer from the data skewness and pipeline bubble issues.
LLM generation optimizations. Many systems focus on optimizing the LLM generation performance [37, 48, 38, 49, 50], and many of them have been adopted by RLHFuse to accelerate the generation stage. Besides, AlpaServe [51] multiplexes the LLM execution with colocated parallelism to improve the inference throughput under bursty workload. FastServe [27] proposes preemptive scheduling on iteration-level to mitigate the head-of-line blocking caused by long-tailed samples. Splitwise [52] and DistServe [26] split the prefill and decoding phases to avoid interference between them. These works are orthogonal to RLHFuse and can be integrated into RLHFuse.
LLM training optimizations. LLM training has been studied by many research works from various aspects. In particular, GPipe [18] and (Interleaved) 1F1B [53, 54] use pipeline parallelism to improve the training throughput and reduce memory footprint. Chimera [20] combines bi-directional pipeline parallelism to further reduce pipeline bubbles. Alpa [55] automatically generates the optimal parallel strategy to improve training performance. MegaScale [17] provides a detailed experience in building a production LLM training system at a large scale. DistTrain [56] focuses on addressing the model and data heterogeneity in multimodal LLM training. These works are agnostic to RLHF thus do not utilizes the characteristics to optimize the RLHF training stage.
@@ 9 Conclusion
We present RLHFuse, a highly efficient RLHF training framework for LLMs. By carefully revisiting the inherent characteristics and task dependencies, RLHFuse views the RLHF workflow from a finer-grained subtask-level perspective while preserving the original synchronous training semantics. This opens up opportunities for efficient inter- and intra-stage fused execution, greatly mitigating the data skewness and pipeline bubbles in existing systems. The evaluation shows that RLHFuse can significantly increase GPU utilization, boosting training throughput by up to 3.7×3.7$\times$3.7 × compared to current state-of-the-art systems.

### v3

@@ 7.3 Effectiveness of Stage Fusion
Inter-stage fusion. One key parameter in inter-stage fusion is the migration threshold $R_{t}$. As $R_{t}$ increases from zero, it shifts from fully serial execution to more aggressive fused execution. However, if $R_{t}$ is set too large, the long-tail generation instances may be overloaded, extending the original generation time. We measure the execution time of the fused generation and inference stage under different migration ratios ($R_{t}/BS$) as shown in Figure 9. We can see the optimal latency is achieved when the remaining samples account for about 20% of the batch size. Note that $R_{t}$ is related to the output length distribution of the model and this experiment is conducted during the early stage of training. As training progresses, $R_{t}$ requires periodic adjustments to adapt to the distribution change. In practice, such changes are usually not too drastic, so the adjustment frequency remains low.
Settings Latency Speedup Peak Memory
@@
Models PP0 PP1 GBS 1F1B+ Greedy Ours LB Greedy Ours
33B/13B 8 4 8 1.10 1.29 1.38 1.38 1.51 1.0
16 1.06 1.12 1.30 1.30 2.14 1.0
32 1.03 1.10 1.15 1.15 2.75 1.26
8 8 1.07 1.4 1.4 1.4 1.51 1.0
16 1.05 1.15 1.32 1.32 2.00 1.19
32 1.03 1.08 1.17 1.17 2.75 1.31
65B/33B 16 8 16 1.05 1.28 1.48 1.48 1.61 1.0
32 1.03 1.14 1.27 1.27 2.23 1.0
64 1.02 1.12 1.15 1.15 2.88 1.26
16 16 1.03 1.27 1.5 1.5 1.57 1.0
32 1.02 1.16 1.33 1.33 2.00 1.22
64 1.01 1.09 1.16 1.17 2.88 1.47
Table 3: The latency speedup and peak activation memory cost of pipeline schedules found by different algorithms under different models, pipeline stages, and global batch size settings.
Intra-stage fusion. We compare RLHFuse’s simulated annealing algorithm against the greedy approach mentioned in §5.2 and show the latency speedup and peak activation memory usage relative to the serial execution of the two models with 1F1B schedule in Table 3. For the latency, a naive solution (denoted as 1F1B+) that does not employ a fused pipeline involves using a smaller PP size without out-of-memory errors. However, this approach increases the DP size and results in fewer micro-batches allocated to each pipeline, thereby partially negating the benefits brought by shallower pipelines. Consequently, this solution is less effective compared to the fused pipeline schedule. We also include a lower bound (denoted as LB) estimation calculated as follows: we first compute the earliest possible completion time for each stage, which consists of three parts: the earliest possible arrival time of the first task, the total time required to process all the tasks, and the remaining time needed for the subsequent pipeline stages of the final task; we then take the maximum of these times across all stages as the lower bound of the fused pipeline schedule. Note that there may not necessarily exist a schedule that reaches this lower bound, but we can use it as a metric to assess how close our algorithm is to the optimal solution. Under all the settings, our approach outperforms the baseline methods and achieves the theoretical lower bound except for the last case.
As for the memory cost, the serial 1F1B execution serves as the lower bound for the fused schedule. Our method also greatly outperforms the greedy approach and often reaches the lower bound. Even when it does not achieve the lower bound, the extra overhead remains within an acceptable range.
Figure 10: The fused pipeline schedule of 65B Actor and 33B Critic model generated by RLHFuse. The two models have 16 and 8 PP stages respectively and #micro-batches is set equal to PP. Top: the GPU execution timeline. Bottom: the peak activation memory usage timeline. The red dotted horizontal line denotes the memory cost of serial 1F1B execution, which is the lower bound.
@@ 7.4 Deep Dive
In this section, we deep dive into RLHFuse by showing a fused pipeline schedule produced by RLHFuse for the 65B/33B model setting. The upper of Figure 10 shows the GPU execution timeline, with forward and backward tasks in light and dark colors, respectively. With intra-stage fusion, RLHFuse strategically fuses one 65B model of 16 PP stages (blue grids) with two 33B models that each have 8 PP stages (pink grids) to fill the pipeline bubbles of each other in reversed pipeline directions. Ultimately, the execution time of this fused pipeline schedule is the same as individually executing the 65B model using the 1F1B schedule, which means we completely overlap the training of the 33B model, achieving the theoretical lower bound. Meanwhile, as shown at the bottom of Figure 10, the peak activation memory usage also achieves the lower bound of executing the two models in serial with 1F1B schedule.
@@ 8 Related Work
RLHF training systems. As RLHF gradually becomes the primary approach for LLM alignment, many frameworks are specifically designed for RLHF training. OpenRLHF [38] uses Ray [47] to distribute models onto separate GPUs and adopts vLLM [8] to accelerate the generation stage. HybridFlow [5] proposes a hierarchical hybrid programming model for the RLHF dataflow and optimizes the GPU allocation and placement of each RLHF model. ReaLHF [6] proposes to design custom parallel strategies for different RLHF tasks to avoid GPU under-utilization. PUZZLE [7] utilizes a lightweight context-switching algorithm to reduce the task switching overhead. These solutions do not consider subtask-level optimization opportunities, thus suffering from data skewness and pipeline bubble issues.
LLM generation optimizations. Many systems focus on optimizing the LLM generation performance [9, 30, 10, 12, 48], and most of them have been adopted by RLHFuse to accelerate the generation stage. Besides, AlpaServe [27] multiplexes the LLM execution with colocated parallelism to improve the throughput under the bursty workload. FastServe [29] proposes preemptive scheduling on iteration-level to mitigate the head-of-line blocking caused by long-tailed samples. Splitwise [49] and DistServe [28] split the prefill and decoding phases to avoid interference between them. These works are orthogonal to RLHFuse and can be integrated into RLHFuse.
LLM training optimizations. LLM training has been studied by many works from various aspects. In particular, [17, 50, 51, 20] use pipeline parallelism to improve the training throughput and reduce memory footprint. Alpa [26] automatically generates the optimal parallel strategy to improve training performance. MegaScale [16] provides a detailed experience in building a production LLM training system at a large scale. DistTrain [52] focuses on addressing the model and data heterogeneity in multimodal LLM training. These works are agnostic to RLHF thus do not utilizes the characteristics to optimize the RLHF training stage.
@@ 9 Conclusion
We present RLHFuse, an efficient RLHF training system. RLHFuse views the RLHF workflow from a finer-grained subtask-level perspective and opens up opportunities for efficient inter- and intra-stage fused execution, mitigating data skewness and pipeline bubbles in existing systems. The evaluation shows that RLHFuse can significantly increase GPU utilization, boosting training throughput by up to 3.7×3.7$\times$3.7 × compared to existing systems.
Acknowledgments. We thank our shepherd, Peter Pietzuch, and the anonymous reviewers for their valuable feedback. We thank Shihong Deng, Shilei Jiang, and Heng Wang for providing algorithmic and data support in this work. We thank Wei Fu and Yi Wu for their assistance in running the baseline and for their valuable feedback on the paper.